The only network issue that Pui Hin had to report was the Helco power outage on Feb 26 from 10:55am to 1:30pm.
Mac Cooper reported an intermittent problem with the connection to Cambridge. He indicated that the connection would become very slow and that a traceroute showed a different path from normal.
Pui Hin asked Mac to send her a normal traceroute and another from when the problem occurs to help her troubleshoot the problem.
Fiber work for the extension has been completed by Oceanic in Waimea and Hilo. Cisco switches that will go to CFHT and IfA Hilo are on order.
UH-ITS has made progress on the negotiations with Hawaiian Telcom on the upgrade to GE for current 100mbps users from Hilo and Waimea to the summit.
The logistics on how to install the gigabit switched network with the least disruption to the summit network were discussed. It was agreed that the summit switch will first be installed with HP connected to it. If everything tests ok, the observatories will schedule with Pui Hin and Miranda for connecting their gigabit switch to the summit switch on their own time schedule. Observatories were asked to send a list of their active subnets with locations (Hilo, summit or Waimea) to Miranda for routing purpose. Date and time were set for everyone to change the subnet mask of the current submit network so that the gigabit network can use the same address space. IPs for the gigabit network were distributed. Observatories were to let Pui Hin and Miranda know if they need their summit fibers to be tested.
Pui Hin reminded everyone of the two designs that were previously considered: The first one is based on pulling fibers from each observatory building to the IfA building. The second one is based on pulling fibers from each observatory building into the manhole where the current fiber panel is located and replacing the fiber panel with a weather proof splice box. The second design will shorten the fiber lengths for the observatories and provides for future expansion by installing more fibers between the manhole and the IfA building than currently needed. The provision for future expansion is desirable as the conduits into the IfA building are getting congested. Pui Hin asked each of the Hilo observatories to review their requirements and let Miranda know if the count per previous discussions is still correct. Here are the number of fibers requested previously. CSO will continue to patch into Gemini.
Gemini - 6MM/6SM JAC - 12MM SMA - 12MM Subaru - 8MM+4SM or 4MM+4SM depending on the cost
New quotes will be obtained once confirmations from all the observatories are received.
Note: this meeting has been rescheduled for June 21st.